During a 5-day festival, the number of visitors tripled each day. If the festival opened on a Thursday with 345 visitors, what was the attendance on that Sunday? 9315 is the answer I need to know the equation
Answer by stanbon(48558)

During a 5-day festival, the number of visitors tripled each day. If the festival opened on a Thursday with 345 visitors, what was the attendance on that Sunday? 9315 is the answer I need to know the equation
---------------------
Thurs = 345
Fri = 345*3
Sat = 345*3*3
Sun = 345*3*3*3
------------------------
EQUATION:
Attendance = 345*3^x where x is the number of days after the 1st day.
========================
Cheers,
Stan H.